OpenStudy (anonymous):

What is the surface area of the prism? A rectangular prism with length of 7 meters, width of 2 meters and height of 4 meters.

OpenStudy (anonymous):

Area=100 Equation for surface area of a rectangular prism= 2(wl+hl+hw) 2(2(7)+4(7)+4(2)) 2(36+14)=2(50)=100